# Relying on the Unreliable: The Impact of Language Models’ Reluctance to Express Uncertainty

### v1

@@ 4.4 Discussion
The perception that plain statements indicate confidence for humans illustrates the harms in the status quo of how LMs’ reluctance to use epistemic markers. The lack of epistemic markers is perceived not as neutral language but as confident language. Although currently models struggle to correctly use epistemic markers in calibrated ways (§3), efforts to linguistically calibrate models is necessary as the absence of these calibration efforts presents significant harms for human-AI collaboration. For example, LM hallucinations are not only factually incorrect Ji et al. (2023); Maynez et al. (2020); Zhang et al. (2023), but are also interperted as high certainty due to the lack of epistemic markers. One potential design recommendation is to generate weakeners without explicit elicitation and use plain statements only when the model is confident.
@@
Miscalibration in the use of strengtheners resulted in users interpreting weakeners incorrectly as well. Our findings signal that miscalibration in one dimension (e.g., the incorrect use of strengtheners) leads users to distrust other areas of epistemic markers. As a recommendation, researchers measuring the harms of miscalibration must also consider how miscalibration impacts the users’ mental models of the whole system, rather than the perception of just an individual component. This work ties into existing literature on how mental models of AI systems are affected by numerous complexities such as accuracy Bansal et al. (2021a), warmth McKee et al. (2022), and system updates Bansal et al. (2019b).
@@
Our findings signal that mental models of language models are developed early in LM-interactions, potentially resulting in long-term harms, even after models become calibrated later on. This is consistent with work from Dhuliawala et al. (2023) who conducted human interpretations of numerical uncertainties from LMs. Similarly, we find that users can correct a mental model developed from an underconfident model but struggle to do so with overconfident models. Unfortunately, publicly deployed models are overconfident, creating not only reliability harm now, but also potentially creating long-term algorithmic aversion Dietvorst et al. (2015) to future models.
@@ 5 Origin of Model Overconfidence
Our work illustrates that LMs are overconfident and that humans are highly reliant on plain statements and statements with strengtheners; all of which is exacerbated by the fact that models are often incorrect when expressing certainty. Here, we turn to our last motivating question: What is the origin of model confidence and what are potential mitigation solutions moving forward? In this section, we pinpoint LM overconfidence to an artifact of the RLHF process, specifically a bias from human annotators against uncertainty.
5.1 Where Does the Overconfidence of Models Begin?
Current state-of-the-art models are trained using a number of techniques to support human-AI collaboration through natural language. Starting with a pretrained base model, one of the most popular techniques is to use supervised fine-tune (SFT) using human instructions. The model is then trained with reinforcement learning with human feedback (RLHF), where a reward model is learned through human preferences of pairwise comparisons of texts.
@@ 5.2 Methods
We compare three models from the GPT3 family, davinci, text-davinci-002, and text-davinci-003 which are base, supervised fine-tuned, and RLHF models respectively. [^9] Models were accessed June - November 2023. We then compare LLaMA-2 models base models with their SFT+RLHF counterparts.
@@
We then directly probe a open-sourced reward model trained on human feedback datasets and assess their scoring. We prompt the model with a question-response pair where the question is "What is the capital of X?" and the response is an epistemic marker like "I think it’s". We test the reward model on 183 question answer pairs across a subset of 30 commonly occurring templates generated by LMs in §3 and compare the model scores with human judgements from §4.
@@
Lastly, we examine the datasets that were used to train the open-sourced reward model. Specifically, we examine the datasets: OpenAI’s “WebGPT comparison” and “Summarize with Feedback”, Dahoa’s “Synthetic Instruct GPT Pairwise” dataset, and Anthropic’s “Helpful and Harmless” dataset. [^10] Reward model and link to datasets: https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2 We then measure how often strengtheners and weakeners are preferred by human annotators in these datasets.
@@ 5.3 Findings
Figure 6: Base models vs RLHF models in their generation of strengtheners and weakeners. In base models, we see a preference for weakeners but the trend reverses among RLHF models.
@@
Reward modeling prefers plain statements with an average score of 4.03, followed by strengtheners with a score of 0.82 (Table 3). However, there is a strong penalty applied to weakeners, with the average rewards score of -1.86. Lower scores in reward modeling would in turn influence how LMs generate natural language, leading to a bias in avoiding the language of uncertainty at generation time.
Marker Human Reward
plain 0.863 4.029
strengthener 0.894 0.818
weakener 0.095 -1.855
Table 3: Comparison between human certainty scores and reward scores by OpenAssistant’s reward model.
@@
We annotate texts in each datasets as containing strengtheners and weakeners and measure how these rates vary across the chosen and rejected texts as annotated by human judges. Given the prevalence of model overconfidence, we hypothesized that there may be a preference for strengtheners in chosen texts. We find that this is not the case. In fact, there is a slight, but significant preference for strengtheners in rejected texts (2.95% vs 2.72%).11111195% CI calculated using bootstrap sampling. In a pairwise comparison, we see that plain text is actually slightly preferred over strengthened texts (chosen 9% more often). This shows that there is not a human bias for strengtheners. However, we do find that weakeners appears significantly more often among rejected texts (5.02%) compared to chosen texts (4.47%). In a pairwise comparison, weakeners are chosen 8% less often than strengtheners and 9% less often than plain texts. This highlights that annotators don’t have a bias for certainty, rather there is a bias against weakeners, matching the results seen from the reward model experiments.
@@ 5.4 Discussion
Uncovering Unknowns in RLHF
Our investigation into RLHF datasets also reveals that humans have implicit biases towards other dimensions of language which may not be known in the annotation phase. The artifact of humans having a bias against uncertainty adds to the list of implicit biases in annotations Gururangan et al. (2018) (e.g., text length Saito et al. (2023), toxicity detection Sap et al. (2022), political leanings Santurkar et al. (2023)). However, the human bias against uncertain language is particularly harmful as it causes RLHF models to be reluctant in their generation uncertainty, negatively impacting human over-reliance on LMs. Simple interventions such as swapping labels on annotated datasets could potentially mitigate some overconfidence harms but could also risk introducing new, yet to be known biases. However, RLHF processes remain significantly understudied due to closed processes and lack of access to datasets and documentation. Investigations are needed in RLHF datasets and that annotation process to fully understand the potential implicit biases that are introduced through human preference annotations.
@@
As LMs evolve towards generating more nuanced language, such as uncertainties, a shift in design is prudent. The work of Hollan and Stornetta (1992) discusses the need to design technology that goes beyond simply mimicking that of the real world. We could apply this design thinking towards designing natural language as an interface: when it comes to expressions of uncertainty, mimicking humans (or mimicking what humans prefer) might not be the end goal. Instead, we could design LMs to verbalize uncertainty in ways that would increase cognitive engagement and lower human overreliance on imperfect models Buçinca et al. (2021). Although humans may have biases in their interpretations and preferences for expressions of uncertainty, we can design LM interfaces to account for or counteract these potential harms.
@@ 7 Limitations
Our study focused exclusively on how language models generate epistemic markers in the English language Bender (2019). Humans greatly differ in their use of hedges and strengtheners across languages, contexts, and cultures Itani (1995); Lauwereyns (2002); Yagız and Demir (2014); Nguyen Thi Thuy (2018); Mur-Duenas (2021), future studies could consider how non-English language models differ in their use of strengtheners and weakeners.
@@ Acknowledgements
Thank you so much to Myra Cheng, Kristina Gligorić, Dan Jurafsky, Abhilasha Ravichander, Tal August, Luca Soldaini and the AI2 Mosaic team for their helpful feedback and advice! Kaitlyn Zhou is supported by the Stanford Graduate Fellowship.
@@ Appendix C Insensitivity to Increases in Temperature
Template Rely %
I can’t answer this question with certainty, … 0.0%
I am not confident, maybe it’s… 0.0%
I am not familiar, maybe it’s… 0.0%
I’m not entirely certain, maybe it’s… 0.0%
I’m not completely sure, maybe it’s… 0.0%
My answer is not definitive, maybe it’s… 0.0%
I have uncertainties about the question, mayb… 0.0%
It’s impossible to say with certainty, maybe … 0.0%
With have some uncertainties it’s… 0.0%
I’m cannot be completely certain, maybe it’s… 0.0%
I’m confused, maybe it’s… 0.0%
I’m not even sure, maybe it’s… 4.0%
I’m not sure, maybe it’s… 4.0%
I’m not sure which of these is correct, maybe… 4.0%
I can’t guarantee, maybe it’s… 4.0%
I am not sure, maybe it’s… 4.0%
I am hesitating, maybe it’s… 4.0%
I cannot confidently say, maybe it’s… 4.0%
I don’t know, maybe it’s… 4.0%
I cannot provide a definitive answer, maybe i… 4.0%
It’s not entirely clear, maybe it’s… 4.0%
I cannot ensure that my answer is entirely co… 4.0%
I may not be entirely accurate, maybe it’s… 4.0%
I may not be entirely correct, maybe it’s… 4.0%
It’s impossible to say for sure, maybe it’s… 4.0%
I’m not 100% sure, maybe it’s… 4.0%
I’m not completely certain, maybe it’s… 4.0%
I am unsure, maybe it’s… 8.0%
I’m not entirely sure, maybe it’s… 8.0%
I am not confident, maybe it’s… 8.0%
It’s hard to be absolutely certain, maybe it’… 8.0%
I’m not absolutely certain, maybe it’s… 8.0%
I’m not 100% confident, maybe it’s… 8.0%
I’m not 100% certain, maybe it’s… 8.0%
It could be… 8.0%
It is not clear, maybe it’s… 8.0%
I’m guessing it’s… 8.0%
I cannot guarantee, maybe it’s… 8.0%
It’s difficult to say, maybe it’s… 8.0%
I am hesitant, maybe it’s… 12.0%
Maybe it’s… 16.0%
While there is some uncertainty, I would gues… 16.0%
I’m not confident, maybe it’s… 20.0%
It is possible it’s… 20.0%
It is probable it’s… 24.0%
I would lean it’s… 32.0%
I’m somewhat confident it’s… 36.0%
I think it’s… 44.0%
It’s more likely it’s… 48.0%
@@
Template Rely %
It seems likely it’s… 52.0%
I’m pretty sure it’s… 52.0%
I would say it’s… 52.0%
It’s very likely it’s… 56.0%
I believe it’s… 64.0%
It’s fairly accurate it’s… 64.0%
I’m fairly certain it’s… 64.0%
It is likely it’s… 68.0%
I would answer it’s… 68.0%
I’m fairly sure it’s… 68.0%
I’m fairly confident it’s… 72.0%
With fair degree of confidence it’s… 72.0%
It’s quite likely it’s… 76.0%
It’s highly likely it’s… 76.9%
It’s extremely likely it’s… 80.0%
I’m pretty certain it’s… 80.0%
I am sure it’s… 84.0%
I’m pretty confident it’s… 84.0%
I’m quite confident it’s… 84.0%
It’s most likely it’s… 84.0%
I feel most confident it’s… 84.0%
Undoubtedly it’s… 84.0%
Without a doubt it’s… 88.0%
I’m sure it’s… 88.0%
I can confidently say it’s… 88.0%
With pretty high certainty it’s… 92.0%
With strong degree of certainty it’s… 92.0%
I’m entirely sure it’s… 92.0%
With high degree of confidence it’s… 92.0%
With high certainty it’s… 92.0%
I’m completely sure it’s… 92.0%
I’m confident it’s… 92.0%
I’m quite sure it’s… 92.0%
It is certain it’s… 92.0%
It’s definitely… 96.0%
I’m entirely confident it’s… 96.0%
I’m very certain it’s… 96.0%
With completely certain it’s… 96.0%
With utmost certainty it’s… 96.0%
With complete certainty it’s… 96.0%
With high degree of certainty it’s… 96.0%
I’m highly confident it’s… 96.0%
I’m quite certain it’s… 96.0%
I’m very confident it’s… 96.0%
With great certainty it’s… 96.0%
I am certain it’s… 96.0%
Without a shred of doubt it’s… 96.0%
I’m absolutely confident it’s… 96.0%
I’m 100% certain it’s… 96.0%
I’m extremely confident it’s… 100.0%
I’m extremely certain it’s… 100.0%
I’m completely confident it’s… 100.0%
We can say with certainty it’s… 100.0%
With absolute certainty it’s… 100.0%
With absolutely certain it’s… 100.0%
I know it’s… 100.0%
I am confident it’s… 100.0%
With full certainty it’s… 100.0%
I’m absolutely sure it’s… 100.0%
Table 7: Human Judgements of Templates Based on Reliability (continued)

### v2

@@ 4.4 Discussion
The troubling concern that current models struggle to use epistemic markers in calibrated ways (§3) is underscored by our finding that lack of epistemic markers is perceived as confident language (§4). That is, the absence of calibrated markers presents significant harms for human-AI collaboration. For example, LM hallucinations are not only factually incorrect Ji et al. (2022); Maynez et al. (2020); Zhang et al. (2023), but as our study suggests, they may also be interpreted as high certainty due to the lack of epistemic markers. One potential design recommendation is to generate weakeners without explicit elicitation and use plain statements only when the model is confident.
@@
Miscalibration in the use of strengtheners resulted in users interpreting weakeners incorrectly as well. Our findings signal that miscalibration in one dimensional (e.g., the incorrect use of strengtheners leads users to distrust other uses of epistemic markers). As a recommendation, researchers measuring the harms of miscalibration must also consider how miscalibration impacts the users’ mental models of the whole system, rather than the perception of just an individual component. This work ties into existing literature on how mental models of AI systems are affected by numerous complexities such as accuracy Bansal et al. (2021a), warmth McKee et al. (2024), and system updates Bansal et al. (2019b).
@@
Our findings signal that mental models of language models are developed early in LM-interactions, potentially resulting in long-term harms, even after models become calibrated later on. This corroborates work from Dhuliawala et al. (2023) who conducted human interpretations of numerical uncertainties from LMs. Similarly, we find that users can correct a mental model developed from an underconfident model but struggle to do so with overconfident models. Unfortunately, public models are overconfident, creating not only a reliability harm now, but also potentially creating long-term algorithmic aversion Dietvorst et al. (2015) to future models.
@@ 5 Origin of Model Overconfidence
Our work shows that LMs are overconfident and that humans are highly reliant on plain statements and statements with strengtheners; all of which is exacerbated by models often being incorrect when expressing certainty. Here, we turn to our last motivating question: What is the origin of model confidence and what are potential mitigations? In this section, we pinpoint LM overconfidence to an artifact of the post-training alignment process, specifically a bias from human annotators against uncertainty.
5.1 Where Does LM Overconfidence Begin?
Current state-of-the-art models are trained using a number of techniques to support human-AI collaboration through natural language. Starting with a pretrained base model, one of the most popular techniques is to use supervised fine-tune (SFT) using human instructions. The model is then trained with reinforcement learning with human feedback (RLHF), where a reward model is learned through human preferences of pairwise text comparisons.
@@ 5.2 Methods
We compare three models from the GPT3 family, davinci, text-davinci-002, and text-davinci-003 which are base, supervised fine-tuned, and RLHF models, respectively. [^10] Models were accessed June - November 2023. We then compare LLaMA-2 models base models with their SFT+RLHF counterparts.
@@
We directly probe OpenAssistant’s open-sourced reward model trained on human feedback datasets and assess their scoring. [^11] https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2 We prompt the model with a question-response pair where the question is "What is the capital of X?" and the response is an epistemic marker like "I think it’s". We test the reward model on 183 question answer pairs across a subset of 30 commonly occurring templates generated by LMs in §3 and compare the model scores with human judgements from §4.
@@
Lastly, we examine the datasets that were used to train the open-sourced reward model. Specifically, we examine the datasets: OpenAI’s “WebGPT comparison” and “Summarize with Feedback”, Dahoa’s “Synthetic Instruct GPT Pairwise” dataset, and Anthropic’s “Helpful and Harmless” dataset.11 We then measure how often strengtheners and weakeners are preferred by human annotators in these datasets.
@@ 5.3 Findings
Figure 6: Base models vs. RLHF models in their generation of strengtheners and weakeners. In base models, we see a preference for weakeners but the trend reverses among RLHF models.
@@
Reward modeling prefers plain statements with an average score of 4.03, followed by strengtheners with a score of 0.82. However, there is a strong penalty applied to weakeners, with the average rewards score of -1.86. Lower scores in reward modeling would in turn influence how LMs generate natural language, leading to a bias in avoiding the language of uncertainty at generation time. See Table 3 for a comparison between reward model and human scoring.
@@
We annotate texts in each datasets as containing strengtheners and weakeners and measure how these rates vary across the chosen and rejected texts as annotated by human judges. Given the prevalence of model overconfidence, we hypothesized that there may be a preference for strengtheners in chosen texts. We find that this is not the case. In fact, there is a slight, but significant preference for strengtheners in rejected texts as compared to chosen text (2.95% vs. 2.72%).12121295% CI calculated using bootstrap sampling. In a pairwise comparison, we see that plain text is actually slightly preferred over strengthened texts (chosen 9% more often). This shows that in these datasets, there is not a human bias for strengtheners. Weakeners also appears significantly more often among rejected texts (5.02%) compared to chosen texts (4.47%). In a pairwise comparison, weakeners are preferred 8% less often than strengtheners and 9% less often than plain texts. This highlights that annotators don’t have a bias for certainty, rather there is a bias against weakeners, matching the results seen from the reward model experiments.
@@ 5.4 Discussion
Uncovering Unknowns in Human Preferences Feedback Alignment
Our investigation into RLHF datasets reveals that humans have implicit biases towards other dimensions of language which may not be known in the annotation phase. Although our study focuses RLHF, the process of aligning language models to biased human feedback will remain as issue, regardless of the exact training method (e.g., DPO or SLiC) Zhao et al. (2023); Rafailov et al. (2024). The artifact of humans having a bias against uncertainty adds to the list of implicit biases in annotations Gururangan et al. (2018) (e.g., text length Saito et al. (2023), toxicity detection Sap et al. (2022b), political leanings Santurkar et al. (2023)). However, the human bias against uncertain language is particularly harmful as it causes aligned models to be reluctant in their generation uncertainty, negatively impacting human over-reliance on LMs. Interventions such as swapping labels on annotated datasets could potentially mitigate some overconfidence harms but could also risk introducing new, yet to be known biases. Investigations are needed in human feedback datasets and annotation processes to fully understand the potential implicit biases that are introduced through human preference annotations.
@@
As LMs evolve towards generating more nuanced language, such as uncertainties, a shift in design is prudent. The work of Hollan and Stornetta (1992) discusses the need to design technology that goes beyond simply mimicking that of the real world. We could apply this design thinking towards designing natural language as an interface. Instead of building language models simply based on mimicking human preferences, we could instead design LMs to verbalize uncertainty in ways that would increase cognitive engagement and lower human overreliance Buçinca et al. (2021).
6 Desired Criteria and Mitigating Solutions Moving Forward
Our research highlights the overconfidence of language models, the reliance of humans on generations, and the long-term consequences of miscalibration in human-AI interactions. Here, we propose three criteria and potential implementations to mitigate overconfidence in LMs.
Criteria: Unsolicited Epistemic Markers
Language models should autonomously emit expressions of uncertainty without prompting, akin to human behavior, in order to appropriately convey levels of confidence. This is critical as the lack of epistemic marker emittance is perceived by humans as tacit certainty.
Concretely, this could be addressed in various stages of the RLHF pipeline, including data augmentation, annotator training, or dataset evaluation. RLHF datasets could be augmented to include additional and more representative epistemic markers. Similar studies have been successful in improving safety Ji et al. (2023), modeling different perspectives Dong et al. (2024), and performing moral self-correction Ganguli et al. (2023). Annotators could also receive training Clark et al. (2021) and priming Sap et al. (2019) to be cognizant of potential biases (e.g., biases against uncertainty). Lastly, automatic sensibility checks and data filtering using our heuristics could be used to evaluate the effectiveness of the proceeding interventions, building transparency towards the levels of overconfidence in RLHF datasets.
Criteria #2: Comprehensive Coverage
Language models ought to have the capacity to generate and interpret the full range of epistemic markers. However, current pre-training data is often limited to written internet data, leading to numerous limitations and biases Gordon and Van Durme (2013); Lucy and Gauthier (2017); Sap et al. (2022a).
Concretely, incorporating more diverse data sources would allow models to learn a larger range of epistemic markers. For example, hedges have long been studied in speech in formal and informal settings Aijmer (1986, 2013) such as presidential debates and speeches Al-Rashady (2012); Mansour and Alghazo (2021), TED talks Nuraniwati and Permatasari (2021), and student presentations Muziatun et al. (2021). Leveraging alternative training sources such as podcasts, news transcripts, and peer conversations could broaden the use of expressions of uncertainty by LMs.
Criteria #3: Context-Dependent Calibration
In addition to being internally calibrated (i.e., calibrated between internal probabilities and verbalized certainties), LMs should also be context-dependent in their calibration. If a model were deployed for entertainment purposes, a lower threshold for expressions of certainty could be tolerated; but if the same language model were to be deployed to a mission-critical task, new thresholds based on the context must be determined.
Concretely, this could be accomplished via user calibration or system-level prompts. For example, one could implement an initialization procedure, similar to the procedure shown in §4.2, that allows practitioners to measure how humans would rely on LM-generated expressions in a specific context. The results of this procedure would then inform practitioners and users of how to best to fit the model to the specific needs using few-shot prompting, fine-tuning, or system-level prompts. Future work could include investigate online algorithms that adjust the certainty thresholds based on the human-LM initialize round.
@@ 8 Limitations
Our study focused exclusively on how language models generate epistemic markers in the English language Bender (2019). Humans greatly differ in their use of hedges and strengtheners across languages, contexts, and cultures Itani (1995); Lauwereyns (2002); Yagız and Demir (2014); Nguyen Thi Thuy (2018); Mur-Dueñas (2021), future studies could consider how non-English language models differ in their use of strengtheners and weakeners.
@@
9 Ethics Statement
Our paper is primarily considered with the potential harms and ethical implications that may arise when humans interact with LLMs. Our findings illustrate that LLMs systematically fail to represent model uncertainties, creating a threat that results humans overrelying on incorrect generations. In regards to our human experiments, we followed standard practices such as providing participants with informed consent, paying participants, on average, over \$15 USD/per hour and debriefing participants on the correct answers if they made errors when interacting with our system.
@@ Acknowledgements
Thank you so much to Dan Jurafsky, Yejin Choi, Myra Cheng, Kristina Gligorić, Abhilasha Ravichander, Tal August, Luca Soldaini and the AI2 Mosaic team for their helpful feedback and advice! Thank you so much to the numerous participants who helped us with our pilot and formal human studies. Kaitlyn Zhou is supported by the Stanford Graduate Fellowship. Xiang Ren’s research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006, the Defense Advanced Research Projects Agency with award HR00112220046, and NSF IIS 2048211.
@@ Appendix C Insensitivity to Increases in Temperature
Appendix D Reward Model Compared to Human Scoring of Expressions of Uncertainty
Marker Human Reward
plain 0.863 4.029
strengthener 0.894 0.818
weakener 0.095 -1.855
Table 3: Comparison between human certainty scores and reward scores by OpenAssistant’s reward model. Human scores are calculated as the percentage of time humans relied on an expression in the control setting.
Appendix E Licenses for Scientific Artifacts
All artifacts were used as intended (querying the models and using the datasets for evaluation and analysis). Used artifacts include: OpenAssistant Reward model and Datasets (License: MIT), MMLU Dataset (License: MIT), and LLaMA 2 (LLaMA 2 Community License).
@@ Appendix E Licenses for Scientific Artifacts
Template Rely % Template Rely %
I can’t answer this question with certainty, … 0.0% I believe it’s… 64.0%
I am not confident, maybe it’s… 0.0% It’s fairly accurate it’s… 64.0%
I am not familiar, maybe it’s… 0.0% I’m fairly certain it’s… 64.0%
I’m not entirely certain, maybe it’s… 0.0% It is likely it’s… 68.0%
I’m not completely sure, maybe it’s… 0.0% I would answer it’s… 68.0%
My answer is not definitive, maybe it’s… 0.0% I’m fairly sure it’s… 68.0%
I have uncertainties about the question, mayb… 0.0% I’m fairly confident it’s… 72.0%
It’s impossible to say with certainty, maybe … 0.0% With fair degree of confidence it’s… 72.0%
With have some uncertainties it’s… 0.0% It’s quite likely it’s… 76.0%
I’m cannot be completely certain, maybe it’s… 0.0% It’s highly likely it’s… 76.9%
I’m confused, maybe it’s… 0.0% It’s extremely likely it’s… 80.0%
I’m not even sure, maybe it’s… 4.0% I’m pretty certain it’s… 80.0%
I’m not sure, maybe it’s… 4.0% I am sure it’s… 84.0%
I’m not sure which of these is correct, maybe… 4.0% I’m pretty confident it’s… 84.0%
I can’t guarantee, maybe it’s… 4.0% I’m quite confident it’s… 84.0%
I am not sure, maybe it’s… 4.0% It’s most likely it’s… 84.0%
I am hesitating, maybe it’s… 4.0% I feel most confident it’s… 84.0%
I cannot confidently say, maybe it’s… 4.0% Undoubtedly it’s… 84.0%
I don’t know, maybe it’s… 4.0% Without a doubt it’s… 88.0%
I cannot provide a definitive answer, maybe i… 4.0% I’m sure it’s… 88.0%
It’s not entirely clear, maybe it’s… 4.0% I can confidently say it’s… 88.0%
I cannot ensure that my answer is entirely co… 4.0% With pretty high certainty it’s… 92.0%
I may not be entirely accurate, maybe it’s… 4.0% With strong degree of certainty it’s… 92.0%
I may not be entirely correct, maybe it’s… 4.0% I’m entirely sure it’s… 92.0%
It’s impossible to say for sure, maybe it’s… 4.0% With high degree of confidence it’s… 92.0%
I’m not 100% sure, maybe it’s… 4.0% With high certainty it’s… 92.0%
I’m not completely certain, maybe it’s… 4.0% I’m completely sure it’s… 92.0%
I am unsure, maybe it’s… 8.0% I’m confident it’s… 92.0%
I’m not entirely sure, maybe it’s… 8.0% I’m quite sure it’s… 92.0%
I am not confident, maybe it’s… 8.0% It is certain it’s… 92.0%
It’s hard to be absolutely certain, maybe it’… 8.0% It’s definitely… 96.0%
I’m not absolutely certain, maybe it’s… 8.0% I’m entirely confident it’s… 96.0%
I’m not 100% confident, maybe it’s… 8.0% I’m very certain it’s… 96.0%
I’m not 100% certain, maybe it’s… 8.0% With completely certain it’s… 96.0%
It could be… 8.0% With utmost certainty it’s… 96.0%
It is not clear, maybe it’s… 8.0% With complete certainty it’s… 96.0%
I’m guessing it’s… 8.0% With high degree of certainty it’s… 96.0%
I cannot guarantee, maybe it’s… 8.0% I’m highly confident it’s… 96.0%
It’s difficult to say, maybe it’s… 8.0% I’m quite certain it’s… 96.0%
I am hesitant, maybe it’s… 12.0% I’m very confident it’s… 96.0%
Maybe it’s… 16.0% With great certainty it’s… 96.0%
While there is some uncertainty, I would gues… 16.0% I am certain it’s… 96.0%
I’m not confident, maybe it’s… 20.0% Without a shred of doubt it’s… 96.0%
It is possible it’s… 20.0% I’m absolutely confident it’s… 96.0%
It is probable it’s… 24.0% I’m 100% certain it’s… 96.0%
I would lean it’s… 32.0% I’m extremely confident it’s… 100.0%
I’m somewhat confident it’s… 36.0% I’m extremely certain it’s… 100.0%
I think it’s… 44.0% I’m completely confident it’s… 100.0%
It’s more likely it’s… 48.0% We can say with certainty it’s… 100.0%
It seems likely it’s… 52.0% With absolute certainty it’s… 100.0%
I’m pretty sure it’s… 52.0% With absolutely certain it’s… 100.0%
I would say it’s… 52.0% I know it’s… 100.0%
It’s very likely it’s… 56.0% I am confident it’s… 100.0%
With full certainty it’s… 100.0%